If 30.6 calories are required to vaporize 25g of water, what is the heat of vaporization (in cal/mole) of that substance?

1
Expert's answer
2021-12-17T18:15:40-0500

"25g" of water contains


"\nu=\dfrac{25g}{18g/mole} = \dfrac{25}{18}mole"

mole of substance. Thus, the heat of vaporization is the following:


"h = \dfrac{30.6cal}{25/18mole} \approx 22cal/mole"

Answer.  22 cal/mole.